# Gravitational Waves on Kerr Black Holes I: Reconstruction of Linearized Metric Perturbations

### v1

@@ 2.3 Radial modes
The derivation of these modes is presented below in Sec. 2.3. Their behavior near the horizon is best described using the tortoise coordinate (2.2), which remains regular there:
$\hat{R}_{\omega\ell m}^{(s)\,{\rm in}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}\mathopen{}\mathclose{{}\left(r_{+}-r_{-}}\right)^{-s}\mathopen{}\mathclose{{}\left(r-r_{+}}\right)^{-ic_{+}k-s}\approx\Delta^{-s}e^{-ikr_{*}},$ (2.20a)
$\hat{R}_{\omega\ell m}^{(s)\,{\rm out}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}\mathopen{}\mathclose{{}\left(r-r_{+}}\right)^{ic_{+}k}\approx e^{ikr_{*}}.$ (2.20b)
@@ 2.4 Teukolsky–Starobinsky constants
which take the explicit form15
@@ 2.7 Metric reconstruction in ingoing radiation gauge
$h_{nm}^{(+)}$ $=\overline{B}^{\rm IRG}H_{\omega\ell m}^{nm},$ (2.50c)
$h_{nm}^{(-)}$ $=\overline{A}^{\rm IRG}H_{-\overline{\omega},\ell,-m}^{nm},$ (2.50d)
@@
$h_{mm}^{(+)}$ $=\overline{B}^{\rm IRG}H_{\omega\ell m}^{mm},$ (2.50f)
$h_{mm}^{(-)}$ $=\overline{A}^{\rm IRG}H_{-\overline{\omega},\ell,-m}^{mm},$ (2.50g)
@@
$H_{\omega\ell m}^{nm}$ $\equiv-\frac{\epsilon_{g}}{2\sqrt{2}\zeta}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}\mathscr{L}_{2}+\frac{a^{2}\sin{2\theta}}{\Sigma}\mathscr{D}_{0}-\frac{2r}{\Sigma}\mathscr{L}_{2}}\right)\overline{\hat{R}}_{-\overline{\omega},\ell,-m}^{(-2)}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)},$ (2.51b)
$H_{\omega\ell m}^{mm}$ $\equiv-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}-\frac{2}{\overline{\zeta}}}\right)\mathscr{D}_{0}\overline{\hat{R}}_{-\overline{\omega},\ell,-m}^{(-2)}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)}.$ (2.51c)
@@ 2.8 Metric reconstruction in outgoing radiation gauge
The Boyer-Lindquist components of the real metric perturbation $h_{\mu\nu}^{\rm ORG}$ are then given by18
@@
$h_{lm}^{(+)}$ $=A^{\rm ORG}H_{\omega\ell m}^{ll},$ (2.57c)
$h_{lm}^{(-)}$ $=B^{\rm ORG}H_{-\overline{\omega},\ell,-m}^{ll},$ (2.57d)
$h_{l\overline{m}}^{(\pm)}$ $=h_{lm}^{(\mp)},$ (2.57e)
@@ 3.2 Application to the Kerr black hole
The operators (3.4) and (3.4) are only well-defined when acting on quantities $\Psi$ with definite GHP weights. If ${\mathcal{S}_{0}^{\dagger}}_{\mu\nu}$ acts on the IRG Hertz potential $\Psi_{\rm H}$ (or any quantity with weight $\mathopen{}\mathclose{{}\left\{-4,0}\right\}$), or if ${\mathcal{S}_{4}^{\dagger}}_{\mu\nu}$ acts on the ORG Hertz potential $\Psi_{\rm H}^{\prime}$ (or any quantity weights $\mathopen{}\mathclose{{}\left\{4,0}\right\}$), then we have
@@ 3.4 Metric reconstruction
Finally, we can put everything together to derive the metric reconstruction formulas presented in Secs. 2.7 and 2.8. The metric perturbations $h_{\mu\nu}$ in IRG and ORG are given by Eqs. (1.18) and (1.20), respectively, with the operators ${S_{0}^{\dagger}}_{\mu\nu}$ and ${S_{4}^{\dagger}}_{\mu\nu}$ taking the forms given in Eqs. (3.10). We apply the operators in this form to the mode expansions (3.11) of the IRG and ORG Hertz potentials. Since each operator has three terms, we see that the metric components $h_{\mu\nu}$ can all be expressed in terms of three functions, which we called $H_{\omega\ell m}^{ab}(r,\theta)$ in Secs. 2.7 and 2.8. The computation of these functions is essentially the main nontrivial step in metric reconstruction.18
@@
Each of these terms takes the form of a second-order differential operator applied to one of the modes of the Hertz potential. We can now replace all the operators $l$, $\frac{\Sigma}{\Delta}n$, $\mathcal{L}_{n}$, and ¯$\overline{\mathcal{L}}_{n}$ by their “mode versions” (2.33), which is how they act on modes behaving like $e^{-i\omega t+im\phi}$. If instead, they act on complex-conjugated modes behaving like $e^{i\overline{\omega}t-im\phi}$, then one must replace $(\omega,m)\to(-\overline{\omega},-m)$ in the operators (2.31) and (2.32), which amounts to sending
$\mathscr{D}_{0}\to\overline{\mathscr{D}}_{0},\qquad\mathscr{D}_{0}^{\dagger}\to\overline{\mathscr{D}}_{0}^{\dagger},\qquad\mathscr{L}_{0}\to\overline{\mathscr{L}}^{\dagger}_{0},\qquad\mathscr{L}_{0}^{\dagger}\to\overline{\mathscr{L}}_{0}.$ (3.16)
@@
$h_{nm}^{(+)}=\overline{B}^{\rm IRG}H_{\omega\ell m}^{nm},\qquad h_{nm}^{(-)}=\overline{A}^{\rm IRG}H_{-\overline{\omega},\ell,-m}^{nm},$ (3.18)
@@
$H_{\omega\ell m}^{nm}\equiv-\frac{\epsilon_{g}}{2\sqrt{2}\zeta}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}\mathscr{L}_{2}+\frac{a^{2}\sin{2\theta}}{\Sigma}\mathscr{D}_{0}-\frac{2r}{\Sigma}\mathscr{L}_{2}}\right)\overline{\hat{R}}_{-\overline{\omega},\ell,-m}^{(-2)}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)}.$ (3.19)
@@ A.2 Linearized Weyl tensor
$C_{\mu\nu\rho\sigma}^{(1)}$ $=\nabla_{\nu}\nabla_{[\rho}h_{\sigma]\mu}-\nabla_{\mu}\nabla_{[\rho}h_{\sigma]\nu}-{h^{\tau}}_{[\rho}R_{\sigma]\tau\mu\nu}-\frac{1}{3}\mathopen{}\mathclose{{}\left(\nabla^{\mu}\nabla^{\nu}h_{\mu\nu}-\nabla^{2}h}\right)\mathopen{}\mathclose{{}\left(g_{\mu\rho}g_{\nu\sigma}-g_{\mu\sigma}g_{\nu\rho}}\right).$ (A.8)
@@ B.1 Metric reconstruction in ingoing coordinates
with $\mathcal{M}_{\pm}$, $\mathcal{M}_{\pm}$, and $\mathcal{N}_{\pm}$ as defined in Eqs. (2.48) and (2.55). Now, we only need to specify the tetrad components (2.5) in these new coordinates. Since the projections $h_{ab}=a^{\mu}b^{\nu}h_{\mu\nu}$ are spacetime scalars, they transform simply by substitution of (B.1). By (B.3), the result is just
@@ B.2 Metric reconstruction in outgoing coordinates
with $\mathcal{M}_{\pm}$, $\mathcal{M}_{\pm}$, and $\mathcal{N}_{\pm}$ as defined in Eqs. (2.48) and (2.55). Now, we only need to specify the tetrad components (2.5) in these new coordinates. Since the projections $h_{ab}=a^{\mu}b^{\nu}h_{\mu\nu}$ are spacetime scalars, they transform simply by substitution of (B.7). By (B.9), the result is just
@@ Appendix C Eliminating derivatives of the mode functions
$\hat{\mathscr{C}}_{\omega\ell m}^{\prime}\hat{R}_{\omega\ell m}^{(-2)}$ $=\mathbb{C}_{\omega\ell m}\Delta\mathscr{D}_{0}^{\dagger}\hat{R}_{\omega\ell m}^{(+2)}+\mathbb{D}_{\omega\ell m}\hat{R}_{\omega\ell m}^{(+2)},$ (C.1b)
@@
$\hat{D}_{\omega\ell m}S_{\omega\ell m}^{(-2)}$ $=\mathbb{Z}_{\omega\ell m}\mathscr{L}_{2}\hat{S}_{\omega\ell m}^{(+2)}+\mathbb{W}_{\omega\ell m}\hat{S}_{\omega\ell m}^{(+2)},$ (C.1d)
where we (yet again) introduced some new coefficients15
@@
$\mathbb{B}_{\omega\ell m}$ $=\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+6-2i\omega r}\right)\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+10+6i\omega r}\right)-12i\omega\mathopen{}\mathclose{{}\left(iK+r-M}\right)}\right]\Delta$
@@
$\mathbb{C}_{\omega\ell m}$ $=-8iK^{3}+K\mathopen{}\mathclose{{}\left\{4i\Delta\lambda_{\omega\ell m}^{(+2)}+8i\mathopen{}\mathclose{{}\left[3\Delta-\mathopen{}\mathclose{{}\left(r-M}\right)^{2}}\right]}\right\}-8i\omega\Delta\mathopen{}\mathclose{{}\left[\Delta-r\mathopen{}\mathclose{{}\left(r-M}\right)}\right],$ (C.2c)
$\mathbb{D}_{\omega\ell m}$ $=\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+2i\omega r+4}\right)\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}-6i\omega r+6}\right)-4i\omega\mathopen{}\mathclose{{}\left(9r-5M}\right)}\right]\Delta^{2}$
@@
$\phantom{=}\quad+4i\lambda_{\omega\ell m}^{(+2)}K\mathopen{}\mathclose{{}\left[iK+3\mathopen{}\mathclose{{}\left(r-M}\right)}\right]\Delta,$ (C.2d)
$\mathbb{Z}_{\omega\ell m}$ $=-4\mathopen{}\mathclose{{}\left[Q\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+2\cot^{2}{\theta}-2Q^{2}+6}\right)+2a\omega\csc{\theta}}\right],$ (C.2e)
$\mathbb{W}_{\omega\ell m}$ $=\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}-6a\omega\cos{\theta}+4}\right)\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+2a\omega\cos{\theta}-4Q^{2}+4Q\cot{\theta}+6}\right)$
$\phantom{=}\quad-12a\omega\mathopen{}\mathclose{{}\left(Q\sin{\theta}-\cos{\theta}}\right),$ (C.2f)
$\mathbb{X}_{\omega\ell m}$ $=4\mathopen{}\mathclose{{}\left[Q\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+2\cot^{2}{\theta}-2Q^{2}+6}\right)+2a\omega\csc{\theta}}\right],$ (C.2g)
@@
$\phantom{=}\quad-12a\omega\mathopen{}\mathclose{{}\left(Q\sin{\theta}+\cos{\theta}}\right).$ (C.2h)
@@ D.1 Explicit metric components including second derivatives
$H_{\omega\ell m}^{nm}$ $\equiv-\frac{\epsilon_{g}}{2\sqrt{2}\zeta}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}\mathscr{L}_{2}+\frac{a^{2}\sin{2\theta}}{\Sigma}\mathscr{D}_{0}-\frac{2r}{\Sigma}\mathscr{L}_{2}}\right)\overline{\hat{R}}_{-\overline{\omega},\ell,-m}^{(-2)}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)}$ (D.1c)
$=-\frac{\epsilon_{g}}{2\sqrt{2}\zeta}\mathopen{}\mathclose{{}\left(\mathop{}\!\partial_{r}\mathop{}\!\partial_{\theta}+A^{nm}\mathop{}\!\partial_{r}+B^{nm}\mathop{}\!\partial_{\theta}+C^{nm}}\right)\overline{\hat{R}}_{-\overline{\omega},\ell,-m}^{(-2)}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)},$ (D.1d)
$H_{\omega\ell m}^{mm}$ $\equiv-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}-\frac{2}{\overline{\zeta}}}\right)\mathscr{D}_{0}\overline{\hat{R}}_{-\overline{\omega},\ell,-m}^{(-2)}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)}$ (D.1e)
$=-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(\mathop{}\!\partial_{r}^{2}+A^{mm}\mathop{}\!\partial_{r}+B^{mm}}\right)\overline{\hat{R}}_{-\overline{\omega},\ell,-m}^{(-2)}\overline{\hat{S}}_{-\overline{\omega},\ell,-m}^{(-2)},$ (D.1f)
@@
$A^{nm}$ $=Q+2\cot{\theta}+\frac{a^{2}\sin{2\theta}}{\Sigma},$ (D.2c)
@@
$C^{nm}$ $=-\mathopen{}\mathclose{{}\left(\frac{iK}{\Delta}+\frac{2r}{\Sigma}}\right)\mathopen{}\mathclose{{}\left(Q+2\cot{\theta}}\right)-\frac{iKa^{2}\sin{2\theta}}{\Sigma\Delta},$ (D.2e)
$A^{mm}$ $=-\frac{2iK}{\Delta}-\frac{2}{\overline{\zeta}},$ (D.2f)
$B^{mm}$ $=\frac{iK}{\Delta}\mathopen{}\mathclose{{}\left(\frac{iK+\Delta^{\prime}}{\Delta}+\frac{2}{\overline{\zeta}}}\right)-\frac{2i\omega r}{\Delta}.$ (D.2g)
@@
$=-\frac{\epsilon_{g}}{8}\frac{\zeta^{2}}{\overline{\zeta}}\Delta^{2}\mathopen{}\mathclose{{}\left(\mathop{}\!\partial_{r}^{2}+A^{mm}\mathop{}\!\partial_{r}+B^{mm}}\right)\hat{R}_{\omega\ell m}^{(+2)}\hat{S}_{\omega\ell m}^{(+2)},$ (D.3f)
where the coefficients of the derivatives are
@@
$A^{lm}$ $=\frac{iK+2\Delta^{\prime}}{\Delta}-\frac{2r}{\Sigma},$ (D.4c)
$B^{lm}$ $=Q+2\cot{\theta}+\frac{a^{2}\sin{2\theta}}{\Sigma},$ (D.4d)

### v2

@@ 2.3 Radial modes
The derivation of these modes is presented below in App. G. Their behavior near the horizon is best described using the tortoise coordinate (2.2), which remains regular there:
$\hat{R}_{\omega\ell m}^{(s)\,{\rm in}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}\mathopen{}\mathclose{{}\left(r_{+}-r_{-}}\right)^{-s}\mathopen{}\mathclose{{}\left(r-r_{+}}\right)^{-ic_{+}k-s}\approx\mathopen{}\mathclose{{}\left[e^{ikr_{+}}\frac{\mathopen{}\mathclose{{}\left(r_{+}-r_{-}}\right)^{-ic_{-}k}}{\mathopen{}\mathclose{{}\left(r_{+}+r_{-}}\right)^{2ikM}}}\right]\Delta^{-s}e^{-ikr_{*}},$ (2.20a)
$\hat{R}_{\omega\ell m}^{(s)\,{\rm out}}(r)$ $\stackrel{{\scriptstyle r\to r_{+}}}{{\approx}}\mathopen{}\mathclose{{}\left(r-r_{+}}\right)^{ic_{+}k}\approx\mathopen{}\mathclose{{}\left[e^{-ikr_{+}}\frac{\mathopen{}\mathclose{{}\left(r_{+}-r_{-}}\right)^{ic_{-}k}}{\mathopen{}\mathclose{{}\left(r_{+}+r_{-}}\right)^{-2ikM}}}\right]e^{ikr_{*}}.$ (2.20b)
@@ 2.4 Teukolsky–Starobinsky constants
which take the explicit form18
@@ 2.7 Metric reconstruction in ingoing radiation gauge
$h_{nm}^{(+)}$ $=\overline{B}^{\rm IRG}\overline{H}_{-\overline{\omega},\ell,-m}^{nm},$ (2.50c)
$h_{nm}^{(-)}$ $=\overline{A}^{\rm IRG}\overline{H}_{\omega\ell m}^{nm},$ (2.50d)
@@
$h_{mm}^{(+)}$ $=\overline{B}^{\rm IRG}\overline{H}_{-\overline{\omega},\ell,-m}^{mm},$ (2.50f)
$h_{mm}^{(-)}$ $=\overline{A}^{\rm IRG}\overline{H}_{\omega\ell m}^{mm},$ (2.50g)
@@
$H_{\omega\ell m}^{nm}$ $\equiv-\frac{\epsilon_{g}}{2\sqrt{2}\overline{\zeta}}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}\mathscr{L}^{\dagger}_{2}+\frac{a^{2}\sin{2\theta}}{\Sigma}\mathscr{D}_{0}-\frac{2r}{\Sigma}\mathscr{L}^{\dagger}_{2}}\right)\hat{R}_{\omega\ell m}^{(-2)}\hat{S}_{\omega\ell m}^{(-2)},$ (2.51b)
$H_{\omega\ell m}^{mm}$ $\equiv-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}-\frac{2}{\zeta}}\right)\mathscr{D}_{0}\hat{R}_{\omega\ell m}^{(-2)}\hat{S}_{\omega\ell m}^{(-2)}.$ (2.51c)
@@ 2.8 Metric reconstruction in outgoing radiation gauge
The Boyer-Lindquist components of the real metric perturbation $h_{\mu\nu}^{\rm ORG}$ are then given by21
@@
$h_{lm}^{(+)}$ $=A^{\rm ORG}H_{\omega\ell m}^{lm},$ (2.57c)
$h_{lm}^{(-)}$ $=B^{\rm ORG}H_{-\overline{\omega},\ell,-m}^{lm},$ (2.57d)
$h_{l\overline{m}}^{(\pm)}$ $=\overline{h}_{lm}^{(\mp)},$ (2.57e)
@@ 3.2 Application to the Kerr black hole
The operators (3.4) and (3.4) are only well-defined when acting on quantities $\Psi$ with definite GHP weights. If ${\mathcal{S}_{0}^{\dagger}}_{\mu\nu}$ acts on the IRG Hertz potential $\Psi_{\rm H}$ (or any quantity of weight $\mathopen{}\mathclose{{}\left\{-4,0}\right\}$), or if ${\mathcal{S}_{4}^{\dagger}}_{\mu\nu}$ acts on the ORG Hertz potential $\Psi_{\rm H}^{\prime}$ (or any quantity of weight $\mathopen{}\mathclose{{}\left\{4,0}\right\}$), then we have
@@ 3.4 Metric reconstruction
Finally, we can put everything together to derive the metric reconstruction formulas presented in Secs. 2.7 and 2.8. The metric perturbations $h_{\mu\nu}$ in IRG and ORG are given by Eqs. (1.18) and (1.20), respectively, with the operators ${S_{0}^{\dagger}}_{\mu\nu}$ and ${S_{4}^{\dagger}}_{\mu\nu}$ taking the forms given in Eqs. (3.10). We apply the operators in this form to the mode expansions (3.11) of the IRG and ORG Hertz potentials. Since each operator has three terms, we see that the metric components $h_{\mu\nu}$ can all be expressed in terms of three functions, which we called $H_{\omega\ell m}^{ab}(r,\theta)$ in Secs. 2.7 and 2.8. The computation of these functions is essentially the main nontrivial step in metric reconstruction.21
@@
Each of these terms takes the form of a second-order differential operator applied to one of the modes of the Hertz potential. We can now replace all the operators $l$, $\frac{\Sigma}{\Delta}n$, $\mathcal{L}_{n}$, and ¯$\overline{\mathcal{L}}_{n}$ by their “mode versions” (2.33), which show how they act on modes behaving like $e^{-i\omega t+im\phi}$. If instead, they act on complex-conjugated modes behaving like $e^{i\overline{\omega}t-im\phi}$, then one must replace $(\omega,m)\to(-\overline{\omega},-m)$ in the operators (2.31) and (2.32), which amounts to sending
$\mathscr{D}_{n}\to\overline{\mathscr{D}}_{n},\qquad\mathscr{D}_{n}^{\dagger}\to\overline{\mathscr{D}}_{n}^{\dagger},\qquad\mathscr{L}_{n}\to\overline{\mathscr{L}}^{\dagger}_{n},\qquad\mathscr{L}_{n}^{\dagger}\to\overline{\mathscr{L}}_{n}.$ (3.16)
@@
$h_{nm}^{(+)}=\overline{B}^{\rm IRG}\overline{H}_{-\overline{\omega},\ell,-m}^{nm},\qquad h_{nm}^{(-)}=\overline{A}^{\rm IRG}\overline{H}_{\omega\ell m}^{nm},$ (3.18)
@@
$H_{\omega\ell m}^{nm}\equiv-\frac{\epsilon_{g}}{2\sqrt{2}\overline{\zeta}}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}\mathscr{L}^{\dagger}_{2}+\frac{a^{2}\sin{2\theta}}{\Sigma}\mathscr{D}_{0}-\frac{2r}{\Sigma}\mathscr{L}^{\dagger}_{2}}\right)\hat{R}_{\omega\ell m}^{(-2)}\hat{S}_{\omega\ell m}^{(-2)}.$ (3.19)
@@ A.2 Linearized Weyl tensor
$C_{\mu\nu\rho\sigma}^{(1)}$ $=\nabla_{\nu}\nabla_{[\rho}h_{\sigma]\mu}-\nabla_{\mu}\nabla_{[\rho}h_{\sigma]\nu}-{h^{\tau}}_{[\rho}R_{\sigma]\tau\mu\nu}-\frac{1}{3}\mathopen{}\mathclose{{}\left(\nabla^{\alpha}\nabla^{\beta}h_{\alpha\beta}-\nabla^{2}h}\right)\mathopen{}\mathclose{{}\left(g_{\mu\rho}g_{\nu\sigma}-g_{\mu\sigma}g_{\nu\rho}}\right).$ (A.8)
@@ B.1 Metric reconstruction in ingoing coordinates
with $\mathcal{L}_{\pm}$, $\mathcal{M}_{\pm}$, and $\mathcal{N}_{\pm}$ as defined in Eqs. (2.48) and (2.55). Now, we only need to specify the tetrad components (2.5) in these new coordinates. Since the projections $h_{ab}=a^{\mu}b^{\nu}h_{\mu\nu}$ are spacetime scalars, they transform simply by substitution of (B.1). By (B.3), the result is just
@@ B.2 Metric reconstruction in outgoing coordinates
with $\mathcal{L}_{\pm}$, $\mathcal{M}_{\pm}$, and $\mathcal{N}_{\pm}$ as defined in Eqs. (2.48) and (2.55). Now, we only need to specify the tetrad components (2.5) in these new coordinates. Since the projections $h_{ab}=a^{\mu}b^{\nu}h_{\mu\nu}$ are spacetime scalars, they transform simply by substitution of (B.7). By (B.9), the result is just
@@ Appendix C Eliminating derivatives of the mode functions
$\hat{\mathscr{C}}_{\omega\ell m}^{\prime}\hat{R}_{\omega\ell m}^{(-2)}$ $=-\mathbb{A}_{\omega\ell m}\Delta\mathscr{D}_{0}^{\dagger}\hat{R}_{\omega\ell m}^{(+2)}+\mathbb{C}_{\omega\ell m}\hat{R}_{\omega\ell m}^{(+2)},$ (C.1b)
@@
$\hat{D}_{\omega\ell m}S_{\omega\ell m}^{(-2)}$ $=-\mathbb{X}_{\omega\ell m}\mathscr{L}_{2}\hat{S}_{\omega\ell m}^{(+2)}+\mathbb{Y}_{-\omega,\ell,-m}\hat{S}_{\omega\ell m}^{(+2)},$ (C.1d)
where we (yet again) introduced some new coefficients18
@@
$\mathbb{B}_{\omega\ell m}$ $=\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+6-2i\omega r}\right)\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+4+6i\omega r}\right)-12i\omega\mathopen{}\mathclose{{}\left(iK+r-M}\right)}\right]\Delta$
@@
$\mathbb{C}_{\omega\ell m}$ $=\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+2i\omega r+4}\right)\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}-6i\omega r+6}\right)-4i\omega\mathopen{}\mathclose{{}\left(9r-5M}\right)}\right]\Delta^{2}$
@@
$\phantom{=}\quad+4i\lambda_{\omega\ell m}^{(+2)}K\mathopen{}\mathclose{{}\left[iK+3\mathopen{}\mathclose{{}\left(r-M}\right)}\right]\Delta,$ (C.2c)
$\mathbb{X}_{\omega\ell m}$ $=4\mathopen{}\mathclose{{}\left[Q\mathopen{}\mathclose{{}\left(\lambda_{\omega\ell m}^{(+2)}+2\cot^{2}{\theta}-2Q^{2}+6}\right)+2a\omega\csc{\theta}}\right],$ (C.2d)
@@
$\phantom{=}\quad-12a\omega\mathopen{}\mathclose{{}\left(Q\sin{\theta}+\cos{\theta}}\right).$ (C.2e)
@@ D.1 Explicit metric components including second derivatives
$H_{\omega\ell m}^{nm}$ $\equiv-\frac{\epsilon_{g}}{2\sqrt{2}\overline{\zeta}}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}\mathscr{L}_{2}^{\dagger}+\frac{a^{2}\sin{2\theta}}{\Sigma}\mathscr{D}_{0}-\frac{2r}{\Sigma}\mathscr{L}_{2}}\right)\hat{R}_{\omega\ell m}^{(-2)}\hat{S}_{\omega\ell m}^{(-2)}$ (D.1c)
$=-\frac{\epsilon_{g}}{2\sqrt{2}\zeta}\mathopen{}\mathclose{{}\left(\mathop{}\!\partial_{r}\mathop{}\!\partial_{\theta}+A^{nm}\mathop{}\!\partial_{r}+B^{nm}\mathop{}\!\partial_{\theta}+C^{nm}}\right)\hat{R}_{\omega\ell m}^{(-2)}\hat{S}_{\omega\ell m}^{(-2)},$ (D.1d)
$H_{\omega\ell m}^{mm}$ $\equiv-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}-\frac{2}{\overline{\zeta}}}\right)\mathscr{D}_{0}\hat{R}_{\omega\ell m}^{(-2)}\hat{S}_{\omega\ell m}^{(-2)}$ (D.1e)
$=-\frac{\epsilon_{g}}{2}\mathopen{}\mathclose{{}\left(\mathop{}\!\partial_{r}^{2}+A^{mm}\mathop{}\!\partial_{r}+B^{mm}}\right)\hat{R}_{\omega\ell m}^{(-2)}\hat{S}_{\omega\ell m}^{(-2)},$ (D.1f)
@@
$A^{nm}$ $=-Q+2\cot{\theta}+\frac{a^{2}\sin{2\theta}}{\Sigma},$ (D.2c)
@@
$C^{nm}$ $=\mathopen{}\mathclose{{}\left(\frac{iK}{\Delta}+\frac{2r}{\Sigma}}\right)\mathopen{}\mathclose{{}\left(Q-2\cot{\theta}}\right)-\frac{iKa^{2}\sin{2\theta}}{\Sigma\Delta},$ (D.2e)
$A^{mm}$ $=-2\mathopen{}\mathclose{{}\left(\frac{iK}{\Delta}+\frac{1}{\zeta}}\right),$ (D.2f)
$B^{mm}$ $=\frac{iK}{\Delta}\mathopen{}\mathclose{{}\left(\frac{iK+\Delta^{\prime}}{\Delta}+\frac{2}{\zeta}}\right)-\frac{2i\omega r}{\Delta}.$ (D.2g)
@@
$=-\frac{\epsilon_{g}}{8}\frac{\zeta^{2}}{\overline{\zeta}}\Delta^{2}\mathopen{}\mathclose{{}\left(\mathop{}\!\partial_{r}^{2}+A^{mm}\mathop{}\!\partial_{r}+B^{mm}}\right)\hat{R}_{\omega\ell m}^{(+2)}\hat{S}_{\omega\ell m}^{(+2)}.$ (D.3f)
Here, the coefficients of the derivatives are
@@
$A^{lm}$ $=Q+2\cot{\theta}+\frac{a^{2}\sin{2\theta}}{\Sigma},$ (D.4c)
$B^{lm}$ $=\frac{iK+2\Delta^{\prime}}{\Delta}-\frac{2r}{\Sigma},$ (D.4d)